OpenStudy (anonymous):

rationalize the denominator. √2/3

OpenStudy (anonymous):

Already done?

OpenStudy (mertsj):

3 is a rational number.

OpenStudy (anonymous):

3 is a rational number. HURRAH! =D

OpenStudy (mertsj):

So the denominator does not need to be rationalized. Unless you really meant: \[\sqrt{\frac{2}{3}}\]

OpenStudy (mertsj):

Is that what you meant to write?

OpenStudy (mertsj):

Assuming it is: \[\sqrt{\frac{2}{3}}=\frac{\sqrt{2}}{\sqrt{3}}\times\frac{\sqrt{3}}{\sqrt{3}}=\frac{\sqrt{6}}{3}\]